# HITTER: A HumanoId Table TEnnis Robot via Hierarchical Planning and Learning

### v1

@@ Abstract
Humanoid robots have recently achieved impressive progress in locomotion and whole-body control, yet they remain constrained in tasks that demand rapid interaction with dynamic environments through manipulation. Table tennis exemplifies such a challenge: with ball speeds exceeding 5 m/s, players must perceive, predict, and act within sub-second reaction times, requiring both agility and precision. To address this, we present a hierarchical framework for humanoid table tennis that integrates a model-based planner for ball trajectory prediction and racket target planning with a reinforcement learning–based whole-body controller. The planner determines striking position, velocity and timing, while the controller generates coordinated arm and leg motions that mimic human strikes and maintain stability and agility across consecutive rallies. Moreover, to encourage natural movements, human motion references are incorporated during training. We validate our system on a general-purpose humanoid robot, achieving up to 106 consecutive shots with a human opponent and sustained exchanges against another humanoid. These results demonstrate real-world humanoid table tennis with sub-second reactive control, marking a step toward agile and interactive humanoid behaviors.
@@ I Introduction
To address these challenges, we adopt a hierarchical framework that separates high-level planning from low-level control. At the high level, a model-based planner estimates the ball trajectory and predicts the striking position, velocity, and timing with high precision. At the low level, a reinforcement learning (RL)–based whole-body controller is trained to execute human-like striking motions while tracking the racket targets specified by the planner. This design directly addresses the identified challenges by: (a) having the planner provide fast trajectory prediction and strike planning, thereby offering a stable interface for the controller and improving sample efficiency during training; (b) training the RL controller on consecutive strikes so that it learns agile motions and reliable balance recovery; and (c) incorporating only two human reference motions into training, through which the controller acquires natural and human-like strikes.
@@
  • We validate our approach through extensive real-world experiments, including both humanoid–humanoid and humanoid–human rallies. Our humanoid achieves up to 106 consecutive shots against a human opponent.
@@ II-A Robotic Table Tennis
Robotic table tennis has emerged as a rich testbed for high-speed perception, planning, and control, requiring rapid reactivity to return balls at competitive speeds. Recent progress has combined advances in hardware, control, and learning. For instance, a lightweight, high-torque robotic arm with model predictive control was developed to execute diverse hit styles with precision [7]. Similarly, a model-based approach for hitting velocity control has demonstrated high success rates in returning balls accurately [8]. On the learning side, seminal work on the use of fast learning-based techniques for playing table tennis have been presented by the Gemini Robotics team at DeepMind [9, 10]. In continuation of this work, amateur human-level competitive performance was achieved through a hierarchical architecture with sim-to-real adaptation [11]. Moreover, beyond specialized robotic table tennis devices, general-purpose humanoid robots have been shown to play table tennis using impedance control [12]. However, in prior work, the humanoid was constrained to stand still without agile locomotion, which limited the effective hitting range. In our work, we focus on human-like and agile reaching motions for table tennis hitting, a capability not previously demonstrated on humanoid robots.
@@ IV-B Ball Trajectory Prediction
where $\mathbf{v}$ and $\mathbf{a}$ are estimated from the first and second derivatives of $p(t)$, respectively. Using this model, we apply explicit step-by-step time integration to predict the ball’s future position and velocity, initialized with the estimates from Sec. IV-A. Given a predefined virtual hit plane at $x=-1.37\,\mathrm{m}$, the corresponding hitting time and position can then be computed.
@@ IV-C Racket-ball Interaction
In addition to making contact with the ball, our objective is to successfully return it, which requires determining the racket’s orientation and velocity at the moment of impact. We assume that, at impact, the racket plane is perpendicular to its velocity vector. Unlike approaches that aim to precisely control the landing position, our goal is to ensure a valid return. Thus, we adopt a simplified post-impact flight model and racket–ball interaction model.
During racket–ball contact, we assume a coefficient of restitution $C_{r}$ along the normal to the racket surface and neglect tangential friction. After impact, the ball is assumed to be subject only to gravity during flight. Given the desired landing position on the table $\hat{\mathbf{p}}_{l}$, the desired hitting position $\hat{\mathbf{p}}_{\mathrm{racket}}$, and a predefined flight time from hitting to landing $\Delta t$, the desired outgoing ball velocity $\mathbf{v}_{o}$ is computed as:
@@ V-A Human Motion References
To encourage human-like motions, we incorporate human motion references into policy training. We use two reference swinging motions: forehand and backhand, both generated using the same pipeline. First, we record a video clip of a human performing the swing. A corresponding SMPL [21] motion clip is reconstructed from the video using GVHMR [22], and then re-targeted to the humanoid robot using GMR [23, 16]. The resulting motion clip contains base pose and joint positions at 30 Hz.
@@ V-B3 Asymmetric Actor-Critic
To provide the critic with additional information unavailable to the policy at deployment, we adopt an asymmetric actor–critic framework for training [25] (Table I). For example, to improve reference tracking, we augment the critic’s observations with the robot body poses $T_{\mathcal{B}}$, which facilitate more accurate return estimation. Since several terms of $r_{g}$ are sparse, the episodic return depends on the number of strikes remaining within an episode. Therefore, we additionally provide the critic with the time left in the current episode, $t_{\mathrm{left}}$. Both the actor and critic are implemented as multi-layer perceptrons (MLPs) with three hidden layers of sizes 512, 256, and 128.
@@ VII-C Future Work
Multi-agent training and serving. In our current humanoid–humanoid experiments, both robots were equipped with the same policy without any joint training of different policies. Even under this simple setting, they were able to sustain a table tennis game. Future work could explore explicit multi-agent training frameworks to further improve competitiveness. Another point is that our robots are not yet capable of serving; in both humanoid–human and humanoid–humanoid settings, a human player is required to initiate the rally. Enabling autonomous serving thus represents another important direction for future work.
@@ VIII Conclusions
We have presented a hierarchical system that enables humanoid robots to play table tennis in the real world. By combining a model-based planner for accurate ball trajectory prediction with a reinforcement learning–based whole-body controller, our approach achieves agile, human-like striking motions under sub-second reaction times. Real-world experiments demonstrate high return success rates, natural forehand/backhand strategies, and rallies of up to 106 consecutive shots against human opponents. Furthermore, we show that two humanoid robots can autonomously sustain rallies against each other, highlighting the robustness and generality of our framework. These results advance humanoid control toward more agile, interactive, and human-level behaviors. Eventually we hope to have humanoids play championship calibre table tennis against skilled opponents.

### v2

@@ Abstract
Humanoid robots have recently achieved impressive progress in locomotion and whole-body control, yet they remain constrained in tasks that demand rapid interaction with dynamic environments through manipulation. Table tennis exemplifies such a challenge: with ball speeds exceeding 5 m/s, players must perceive, predict, and act within sub-second reaction times, requiring both agility and precision. To address this, we present a hierarchical framework for humanoid table tennis that integrates a model-based planner for ball trajectory prediction and racket target planning with a reinforcement learning-based whole-body controller. The planner determines striking position, velocity and timing, while the controller generates coordinated arm and leg motions that mimic human strikes and maintain stability and agility across consecutive rallies. Moreover, to encourage natural movements, human motion references are incorporated during training. We validate our system on a general-purpose humanoid robot, achieving up to 106 consecutive shots with a human opponent and sustained exchanges against another humanoid. These results demonstrate real-world humanoid table tennis with sub-second reactive control, marking a step toward agile and interactive humanoid behaviors.
@@ I Introduction
To address these challenges, we adopt a hierarchical framework that separates high-level planning from low-level control. At the high level, a model-based planner estimates the ball trajectory and predicts the striking position, velocity, and timing with high precision. At the low level, a reinforcement learning (RL)-based whole-body controller is trained to execute human-like striking motions while tracking the racket targets specified by the planner. This design directly addresses the identified challenges by: (a) having the planner provide fast trajectory prediction and strike planning, thereby offering a stable interface for the controller and improving sample efficiency during training; (b) training the RL controller on consecutive strikes so that it learns agile motions and reliable balance recovery; and (c) incorporating only two human reference motions into training, through which the controller acquires natural and human-like strikes.
@@
  • We validate our approach through extensive real-world experiments, including both humanoid-humanoid and humanoid-human rallies. Our humanoid achieves up to 106 consecutive shots against a human opponent.
@@ II-A Robotic Table Tennis
Robotic table tennis has emerged as a rich testbed for high-speed perception, planning, and control, requiring rapid reactivity to return balls at competitive speeds. In our own work, we were very much inspired by the biologically-inspired approach to motion generation in [7, 8]. In this work the research team led by Jan Peters fits a template trajectory while minimizing the robot’s deviations from a fixed “comfort posture”: Human demonstration via kinesthetic teach-in was used in [9] to learn a general mixture of motions policy from motion primitives. The interest in responding to adversaries with opponent modeling was presented in [10] with anticipatory action selection [11], and the design of the Intention-driven Dynamics Model [12] and [13]. The work in [14] applied inverse reinforcement learning to identify strategic elements of play from demonstrations between different players. Our own work has analyzed detailed play from hours of recorded video [15] to determine intent. In Section VII, we briefly discuss integrating this into our humanoid going forward.
Recent progress has combined advances in hardware, control, and learning. For instance, a lightweight, high-torque robotic arm with model predictive control was developed to execute diverse hit styles with precision [16]. Similarly, a model-based approach for hitting velocity control has demonstrated high success rates in returning balls accurately [17]. On the learning side, seminal work on the use of fast learning-based techniques for playing table tennis have been presented by the Gemini Robotics team at DeepMind [18, 19]. In continuation of this work, amateur human-level competitive performance was achieved through a hierarchical architecture with sim-to-real adaptation [20]. In contrast to large-scale training efforts, [21] showed that reinforcement learning can learning to rally from scratch in under 200 trials. Moreover, beyond specialized robotic table tennis devices, general-purpose humanoid robots have been shown to play table tennis using impedance control [22]. However, in prior work, the humanoid was constrained to stand still without agile locomotion, which limited the effective hitting range. In our work, we focus on human-like and agile reaching motions for table tennis hitting, a capability not previously demonstrated on humanoid robots.
@@ IV-B Ball Trajectory Prediction
where $\mathbf{v}$ and $\mathbf{a}$ are estimated from the first and second derivatives of $p(t)$, respectively. Using this model, we apply explicit step-by-step time integration to predict the ball’s future position and velocity, initialized with the estimates from Sec. IV-A. Given a predefined virtual hit plane [7, 8, 29] at $x=-1.37\,\mathrm{m}$, the corresponding hitting time and position can then be computed.
@@ IV-C Racket-ball Interaction
In addition to making contact with the ball, our objective is to successfully return it, which requires determining the racket’s orientation and velocity at the moment of impact. We assume that, at impact, the racket plane is perpendicular to its velocity vector. Unlike approaches that aim to precisely control the landing position, our goal is to ensure a valid return. Thus, we adopt a simplified post-impact flight model and racket-ball interaction model.
During racket-ball contact, we assume a coefficient of restitution $C_{r}$ along the normal to the racket surface and neglect tangential friction. After impact, the ball is assumed to be subject only to gravity during flight. Given the desired landing position on the table $\hat{\mathbf{p}}_{l}$, the desired hitting position $\hat{\mathbf{p}}_{\mathrm{racket}}$, and a predefined flight time from hitting to landing $\Delta t$, the desired outgoing ball velocity $\mathbf{v}_{o}$ is computed as:
@@ V-A Human Motion References
We extend the idea of generating instantaneous racket motion close to a “comfort posture” in [7, 8] to generating continuous humanoid motion close to two swinging references: forehand and backhand. Note that while [9] synthesizes a mixture of motor primitives controller from kinesthetic teach-ins, we use video-based demonstrations. First, we record a video clip of a human performing the swing. A corresponding SMPL [32] motion clip is reconstructed from the video using GVHMR [33], and then re-targeted to the humanoid robot using GMR [34, 26]. The resulting motion clip contains base pose and joint positions at 30 Hz.
@@ V-B3 Asymmetric Actor-Critic
To provide the critic with additional information unavailable to the policy at deployment, we adopt an asymmetric actor-critic framework for training [36] (Table I). For example, to improve reference tracking, we augment the critic’s observations with the robot body poses $T_{\mathcal{B}}$, which facilitate more accurate return estimation. Since several terms of $r_{g}$ are sparse, the episodic return depends on the number of strikes remaining within an episode. Therefore, we additionally provide the critic with the time left in the current episode, $t_{\mathrm{left}}$. Both the actor and critic are implemented as multi-layer perceptrons (MLPs) with three hidden layers of sizes 512, 256, and 128.
@@ VII-C Future Work
Multi-agent training and serving. In our current humanoid-humanoid experiments, both robots were equipped with the same policy without any joint training of different policies. Even under this simple setting, they were able to sustain a table tennis game. Future work could explore explicit multi-agent training frameworks to further improve competitiveness. Another point is that our robots are not yet capable of serving; in both humanoid-human and humanoid-humanoid settings, a human player is required to initiate the rally. Enabling autonomous serving thus represents another important direction for future work.
@@ VIII Conclusions
We have presented a hierarchical system that enables humanoid robots to play table tennis in the real world. By combining a model-based planner for accurate ball trajectory prediction with a reinforcement learning-based whole-body controller, our approach achieves agile, human-like striking motions under sub-second reaction times. Real-world experiments demonstrate high return success rates, natural forehand/backhand strategies, and rallies of up to 106 consecutive shots against human opponents. Furthermore, we show that two humanoid robots can autonomously sustain rallies against each other, highlighting the robustness and generality of our framework. These results advance humanoid control toward more agile, interactive, and human-level behaviors. Eventually we hope to have humanoids play championship calibre table tennis against skilled opponents.